OpenStudy (anonymous):

The Fibonacci sequence can be defined recursively as SEE EQUATION IN COMMENTS. What is the 10th term?

OpenStudy (anonymous):

\[a _{1} = 1\] \[a _{2} = 1\] \[a _{n} = a _{n-1} + a _{n - 2} forn \ge 3\]

OpenStudy (amistre64):

depends on the starting terms

OpenStudy (anonymous):

1,1,2,3,5,8,13,21,34,55,89,144,233,377,610,...

OpenStudy (anonymous):

1,1,2,3,5,8,13,21,34,55,...

OpenStudy (amistre64):

THE fibber starts at 1 and 1 tho right?

OpenStudy (anonymous):

what rubin said

OpenStudy (anonymous):

10th term is 55

OpenStudy (amistre64):

ta ta nee chi near tens guns lazarus rosemines mishunbabies mishunbabiess is ow i recall the first 10 factorial lol

OpenStudy (anonymous):

if you want the nth term and have a calculator y ou can compute \[\frac{1}{\sqrt{5}}(\frac{1+\sqrt{5}}{2})^n-\frac{1}{\sqrt{5}}(\frac{1-\sqrt{5}}{2})^n\]

OpenStudy (amistre64):

1,1,2,6,24,120,720,5040,40320,362880,3628800